OpenStudy (anonymous):

A kennel owner has 270 ft of fencing material to be used to divide a rectangular area into equal pens. find the dimensions that would allow ft^2 for each pen. (2 answers)

OpenStudy (anonymous):

Is there a specific square footage that the question is looking for?

OpenStudy (anonymous):

yeah and it can be written 2 wayss

OpenStudy (anonymous):

No k ft^2 for each pen is being provided. Solve the following equations for L and w, length and width, where the each pen is assumed a square, maximum area for a given perimeter, and the pens share a common boundary:\[\left\{2L+3w=270,w\text{=}\left(\frac{1}{2}\right)L\right\}\]\[\left\{L\to \frac{540}{7},w\to \frac{270}{7}\right\}=\{L\to 77.1429,w\to 38.5714\} \]